manufacturer of microcomputer produces three different models. The following table summarizes wholesale process, material cost per unit and labor cost per unit. Annual fixed costs are $ 25 million.



Microcomputer




Model 1


Model 2


Model 3



Wholesale price per unit


$500


$1,000


$1,500



Material cost per unit


$175


$400


$750



Labor cost per unit


$100


$150


$225



Determine a joint total revenue function for sale of the three different microcomputer models?


Determine an annual total cost function for manufacturing the three models.


Determine the profit function for sales of the three models.


1
Expert's answer
2021-12-28T17:27:39-0500
"\text{\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ Microcomputer }""\def\arraystretch{1.5}
 \begin{array}{c:c:c:c}
 & Model\ 1 & Model\ 2 & Model\ 3 \\ \hline
 Wholesale\\ 
price/unit & \$500 & \$1000 & \$1500 \\
 \hdashline
 Material\\ 
cost/unit & \$175 & \$400 & \$750 \\
 \hdashline
 Labor\\ 
cost/unit & \$100 & \$150 & \$225 \\
 \hdashline
\end{array}"

1. A joint total revenue functions for sale of the three different microcomputer models are:


"TR_1 = 500Q_1,"

"TR_2 = 1000Q_2,"

"TR_3 = 1500Q_3."


A joint total revenue function for sale of the three different microcomputer models is


"TR=TR_1+TR_2+TR_3"

"=500Q_1+1000Q_2+1500Q_3"

2. An annual total cost functions for manufacturing the three models:


"TC_1 = 25000000\/3+(175+100)Q_1,"

"TC_2= 25000000\/3+(400+150)Q_2,"

"TC_3= 25000000\/3+(750+225)Q_3."




"TC=TC_1+TC_2+TC_3"

"=25000000+275Q_1+550Q_2+975Q_3"

3. The profit function for sales of the three models is

"TP=TP_1 +TP_2+TP_3"

"=TR-TC"

"=500Q_1+1000Q_2+1500Q_3"

"-(25000000+275Q_1+550Q_2+975Q_3)"

"=225Q_1+450Q_2+525Q_3-25000000"